You need to accumulate $10,000. To do so, you plan to make deposits of $1,000 per year - with the first payment being made a year from today - into a bank account that pays 8.81% annual interest. Your last deposit will be less than $1,000 if less is needed to round out to $10,000. How many years will it take you to reach your $10,000 goal? Round your answer up to the nearest whole number.
year(s)

How large will the last deposit be? Round your answer to the nearest cent.
$  
Answer #1
10000 = 1000*PVIF(8.81,n), where n is the number of years
PVIF(8.81,n) = 10000/1000 = 10
Further, PVIFA for 8.81% and n = 8 = 10.9528
and for n = 7, it is 9.1470
Actual number of years = 7+(10-9.1470)/(10.9528-9.1470) = 7.4724
rounded off to 7 Years.
However, a deposit of less than 1000 has to be made in the 8th year.
In 7 years the FV will be 1000*9.1470 = $9147
The 8th year deposit = 10000-9147*1.0881 = $47